On January 2, 2017​,

FamilyFamily

Clothing Consignments purchased showroom fixtures for 10,000

​cash, expecting the fixtures to remain in service for five years.

Family has depreciated the fixtures on a​ double-declining-balance basis, with zero residual value. On

September 30,

2018

FamilyFamily

sold the fixtures for

$ 5 comma 500$5,500

cash. Record both depreciation expense for

20182018

and sale of the fixtures on

September 30September 30​,

20182018.

Now, record the sale of the fixtures on

September 30September 30​,

20182018.

Date

Accounts and Explanation

Debit

Credit

Sep. 30
Answer #1
Date Accounts and Explanation Debit Credit
Sep-30 Depreciation Expense $ 1,800.00
Accumulated Depreciation $ 1,800.00

Depreciation for 2017:

=10,000 x 2/5 = 4,000

Depreciation for 2018 (through Sept. 30)

(10,000 - 4,000) x 2/5 x 9/12

=1800

Accumulated Depreciation = 4,000 + 1,800 = 5,800

Date Accounts and Explanation Debit Credit
Sep-30 Cash $ 5,500.00
Accumulated Depreciation $ 5,800.00
Fixtures $ 10,000.00
Gain on Sale of assets $    1,300.00